# Detailed nuclear structure calculations for coherent elastic neutrino-nucleus scattering

### v1

@@ V.2 Experimental event counts
To go from the raw scattering rate to the expected number of counts in an experiment we follow the recipes provided by the COHERENT collaboration in [54] and [55]. The total number of detected CEvNS events is found by integrating over the detected energy taking into account the quenching of nuclear recoils, detector resolution effects and the energy dependent detection efficiency. For the cesium-iodide detector this has the form
@@
while for the argon detector the form is
@@
×$\times\int dE_{R}\,P(E_{ee}|E_{R})\frac{dR}{dE_{\nu}},$
where $P\!E$ is the detected photoelectrons, $N_{T}$ is the number of targets in the detector and $\Phi=\frac{\rm{POT}f_{\pi}}{4\pi d^{2}}$ is total number of neutrinos (per flavor) expected at the target [calculated from the total number of protons on target (POT) and the average number of pions produced per POT, $f_{\pi}$]. The values for the experimental parameters are summarized in table 5.
@@
argon 24.4 kg 27.5 m $1.38\times 10^{23}$ 0.09
cesium-iodide 14.6 kg 19.3 m $3.20\times 10^{23}$ 0.0848
@@
The resulting expected counts for the two detectors’ runs are shown in Fig. 4. The uncertainty in the number of counts is determined from two sources: the uncertainty in the cross section (calculated as above), and from the experimental systematic uncertainty (which is dominated by the 10% uncertainty in the total neutrino flux). We take our systematic uncertainty from COHERENT’s analysis (not including their theoretical uncertainties): 13% for argon and 12% for cesium iodide.
@@
Overall we achieved good agreement reproducing COHERENT’s results, with all predictions well within the errors for both COHERENT’s predictions and experimental results. However, without knowledge of their exact theoretical model it is difficult to reproduce the result precisely. In this instance the result obtained using the form factor derived from the electric-charge distribution provides a useful benchmark to compare against. For the cesium-iodide detector our predictions were systematically larger, but shell model results were comparatively lower than from the electric-charge distribution form factors. This is likely due to the slightly larger charge radius predicted from the shell model, which causes coherence to be lost to a greater extent (for a given momentum transfer). For the argon detector our result is systematically lower than COHERENT’s prediction, in contrast to our increased cross section prediction. As with the argon cross sections, the different form factors all produce very similar results because argon-40 is an even-even nucleus.
@@
It is worth noting that the reduction in the theoretical uncertainty achieved for our cross section predictions has essentially no effect on the total uncertainty of our event rate predictions, since they are dominated by the uncertainty in the neutrino flux.
VI Conclusion
@@
In this work we have computed precise flux-averaged cross section predictions for the CEvNS experiments being carried out by the COHERENT collaboration using argon, germanium and cesium-iodide targets. These predictions were made using both Skyme-Hartree-Fock calculations of weak radii, and shell model calculations of nuclear structure factors. The shell model results included sub-leading nuclear effects such as from spin-orbit coupling and axial currents. We used the Skyme-Hartree-Fock calculation to provide a cross-check of the density distribution obtained with the shell model. The results for the $F_{M}$ response showed very good agreement between these two methods.
@@
Our results are in broad agreement with prior work but have substantially reduced the theoretical errors by up to a factor of 10. Future advances on the experimental side will both reduce the uncertainty on the neutrino flux and increase the number of detected CEvNS events, reducing statistical errors. Combined with these future advancements, our results will pave the way for CEvNS to serve as a precision test of the Standard Model.
@@ Appendix A Point-like nucleon distributions
Full results for the point-like distribution of nucleons in each nuclear target are given in the tables below. The SHF results were used to construct Fermi-form factors for the Weak-charge distribution. The shell model results are given for comparison only and are not used for any subsequent calculations. The shell model point-nucleon radii were calculated using
@@
Table 6: Skyrme-Hartree-Fock predictions of root mean square point like proton and neutron radii.
Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm)
40Ar SkM 3.356 3.445
SLy6 3.330 3.412
SV-min 3.321 3.418
UNEDF1 3.296 3.407
Average 3.326(25) 3.420(17)
70Ge SkM 3.949 4.003
SLy6 3.944 3.988
SV-min 3.922 3.980
UNEDF1 3.904 3.975
Average 3.930(21) 3.986(12)
72Ge SkM 3.965 4.053
SLy6 3.959 4.031
SV-min 3.940 4.031
UNEDF1 3.922 4.028
Average 3.946(20) 4.036(12)
73Ge SkM 3.973 4.077
SLy6 3.965 4.054
SV-min 3.947 4.054
UNEDF1 3.931 4.053
Average 3.954(19) 4.059(11)
74Ge SkM 3.980 4.099
SLy6 3.976 4.081
SV-min 3.955 4.077
UNEDF1 3.941 4.078
Average 3.963(18) 4.084(10)
76Ge SkM 4.016 4.163
SLy6 4.004 4.138
SV-min 3.978 4.129
UNEDF1 3.975 4.140
Average 3.993(20) 4.143(14)
127I SkM 4.702 4.836
SLy6 4.698 4.824
SV-min 4.676 4.814
UNEDF1 4.676 4.822
Average 4.688(14) 4.824(9)
133Cs SkM 4.763 4.898
SLy6 4.759 4.885
SV-min 4.739 4.878
UNEDF1 4.740 4.887
Average 4.750(12) 4.887(8)
@@
Appendix B Comparison to nuclear observables
Here we compare experimental nuclear observables to the theoretical shell model values, namely the nuclear energy levels, as well as the magnetic dipole and electric quadrupole moments and transitions. For each of these observables, we provide the values for 40Ar, 70,72,74,76Ge and 133Cs below, whilst the 73Ge and 127I comparisons can be found in [39].
B.1 Argon-40
Figure 5: Experimental energy levels (keV) for 40Ar compared with levels from the SDPF-NR, SDPF-U, SDPF-MU and EPQQM shell model calculations.
Table 8: Magnetic dipole and electric quadrupole moments, and electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions, between low-lying states of 40Ar. Experimental transition values from [56], while moments taken from [57, 58].
@@ B.1 Argon-40
B.2 Germanium-70,72,74,76
Figure 6: Low-lying experimental energy levels (keV) for germanium isotopes compared with levels from shell model calculations using the JUN45 and jj44b interactions.
Table 9: Magnetic dipole and electric quadrupole moments, for low-lying states of 70,72,74,76Ge. Experimental moments taken from [57, 58, 59].
@@ B.2 Germanium-70,72,74,76
Table 10: Electric quadrupole [B(E2)] transitions between low-lying states of 70,72,74,76Ge. Experimental transitions taken from [60, 61, 62, 63].
@@
B.3 Cesium-133
Figure 7: Experimental energy levels (keV) for 133Cs, compared with levels from SN100PN and GCN5082 shell model calculations for two different valence space truncations.
Table 11: Magnetic dipole and electric quadrupole moments for low-lying states of 133Cs. Experimental moments taken from [64, 57, 58]. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
@@ B.3 Cesium-133
$5/2_{1}^{+}$66.40 -67.47 -63.36 -60.36 30(2) [^1] Sign not specified 3.54 3.62 3.36 3.50 +3.45(2)
@@
Table 12: Electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions between low-lying states of 133Cs. Experimental transition values from [65]. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
@@ Appendix C EFT/Nuclear Operators
The nuclear operators employed in Eq. (IV.1) have the form
@@
The one-nucleon matrix elements in Eq. (IV.1) have analytic expressions of the form [15]:
@@ Appendix D Form factors
In this section we plot the form factors for the individual responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ using the shell model calculations performed in this work, where the $\Sigma^{\prime}$ response is only plotted for isotopes with non-zero nuclear ground state spin. This is presented for 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs, highlighting the corresponding COHERENT range of interest for each of these nuclei. The form factor expressions employed to plot these are presented in [39], except for 133Cs which is provided in Appendix D. The KN and SHF form factors were plotted using $(ZF\left(q,r_{\text{rms}}^{p})+NF(q,r_{\text{rms}}^{n})\right)^{2}$, while the shell model counterparts were plotted using $\sum_{N,N^{\prime}=p,n}F^{(N,N^{\prime})}_{X,X}(q^{2})$, where $F^{(N,N^{\prime})}_{X,X}(q^{2})$ is defined in Eq. (40).

### v2

@@ IV.2 Experimental event counts
To go from the raw scattering rate to the expected number of counts in an experiment we follow the recipes provided by the COHERENT collaboration in reference [47] and [48]. The total number of detected CEvNS events is found by integrating over the detected energy taking into account the quenching of nuclear recoils, detector resolution effects and the energy dependent detection efficiency. For the cesium-iodide detector this is:
@@
while for the argon detector it is:
@@
×$\times\int dE_{R}\,P(E_{ee}|E_{R})\frac{dR}{dE_{\nu}}.$
Here $P\!E$ is the detected photoelectrons, $N_{T}$ is the number of targets in the detector and $\Phi$ is total number of neutrinos, per flavor, expected at the detector. The latter is computed as $\Phi=\frac{{\rm{POT}}f_{\pi}}{4\pi d^{2}}$, where POT is the total number of protons on target, $f_{\pi}$ is the average number of pions produced per POT, and $d$ is the distance of the detector from the beam-dump. The values for the experimental parameters are summarized in table 5.
@@
Ar 24.4 kg 27.5 m $1.38\times 10^{23}$ 0.09
CsI 14.6 kg 19.3 m $3.20\times 10^{23}$ 0.0848
@@
The resulting expected counts for the two detectors’ runs are shown in Fig. 4. The uncertainty in the number of counts is determined from two sources: the uncertainty in the cross section (calculated as above), and from the experimental systematic uncertainty (which is dominated by the 10% uncertainty in the total neutrino flux, but also includes detector effects such as efficiency cuts and quenching factors). We take our systematic uncertainty from COHERENT’s analyses (not including their theoretical uncertainties): 13% for argon and 12% for cesium iodide [43, 2].
@@
Overall we find good agreement with COHERENT’s results, with all predictions well within the errors for both COHERENT’s predictions and experimental results. However, without knowledge of their exact theoretical cross section, it is difficult to reproduce their result precisely. In this instance the result obtained using the form factor derived from the electric-charge distribution provides a useful benchmark to compare against. For the cesium-iodide detector our predictions were systematically larger, but shell model results were comparatively lower than from the electric-charge distribution form factors. This is likely due to the slightly larger charge radius predicted from the shell model, which causes coherence to be lost to a greater extent (for a given momentum transfer). For the argon detector our result is systematically lower than COHERENT’s prediction, in contrast to our increased cross section prediction. As with the argon cross sections, the different form factors all produce very similar results because argon-40 is an even-even nucleus.
@@
It is worth noting that the reduction in the theoretical uncertainty achieved for our cross section predictions has essentially no effect on the total uncertainty of our event rate predictions, since they are dominated by the systematic uncertainty in the total neutrino flux.
V Conclusion
@@
In this work we have computed precise flux-averaged cross section predictions for the CEvNS experiments being carried out by the COHERENT collaboration using argon, germanium and cesium-iodide targets. These predictions were made using both Skyme-Hartree-Fock calculations of weak radii, and shell model calculations of nuclear structure factors. These were compared with an empirically derived cross section prediction based on charge-radius measurements. We find that the our calculated weak radii and $F_{M}$ nuclear responses showed reasonable agreement between the two methods and the charge radii data. While the shell model results included additional nuclear effects such as from spin-orbit coupling and axial currents, they were not found to alter the predictions significantly. We observe a small difference between the the SHF and shell model calculations, where the latter gives a slightly smaller cross section for germanium and iodine targets. Given that spin-orbit and axial effects are negligible, we conclude that this difference is entirely driven by the density distribution.
@@
These results are in broad agreement with prior work, but have substantially smaller theoretical errors, even after accounting for the systematic difference between our SHF and shell model results. In the case of CsI, the difference between our SHF and shell model results is a factor of 2 times smaller than COHERENT’s prediction. Further work should be done to understand the origin of this discrepancy and implement more advanced uncertainty estimates. Future advances on the experimental side will both reduce the uncertainty on the neutrino flux and increase the number of detected CEvNS events, reducing statistical errors. Combined with these future advancements, our results will pave the way for CEvNS to serve as a precision test of the Standard Model.
@@ Appendix A Point-like nucleon distributions
Full results for the point-like distribution of nucleons in each nuclear target are given in tables 6 and 7. The SHF results were used to construct Fermi-form factors for the Weak-charge distribution. The shell model results are given for comparison only and are not used for any subsequent calculations. The shell model point-nucleon radii were calculated using
@@
Table 6: Skyrme-Hartree-Fock predictions of root mean square point like proton, neutron, and weak radii.
Nucleus Interaction $R_{p}$ (fm) $R_{n}$ (fm) $R_{w}$ (fm)
40Ar SkM 3.356 3.445 3.539
SLy6 3.330 3.412 3.507
SV-min 3.321 3.418 3.513
UNEDF1 3.296 3.407 3.503
70Ge SkM 3.949 4.003 4.083
SLy6 3.944 3.988 4.068
SV-min 3.922 3.980 4.061
UNEDF1 3.904 3.975 4.057
72Ge SkM 3.965 4.053 4.134
SLy6 3.959 4.031 4.112
SV-min 3.940 4.031 4.113
UNEDF1 3.922 4.028 4.111
73Ge SkM 3.973 4.077 4.159
SLy6 3.965 4.054 4.135
SV-min 3.947 4.054 4.136
UNEDF1 3.931 4.053 4.136
74Ge SkM 3.980 4.099 4.181
SLy6 3.976 4.081 4.163
SV-min 3.955 4.077 4.160
UNEDF1 3.941 4.078 4.161
76Ge SkM 4.016 4.163 4.245
SLy6 4.004 4.138 4.220
SV-min 3.978 4.129 4.212
UNEDF1 3.975 4.140 4.223
127I SkM 4.702 4.836 4.907
SLy6 4.698 4.824 4.895
SV-min 4.676 4.814 4.886
UNEDF1 4.676 4.822 4.894
133Cs SkM 4.763 4.898 4.969
SLy6 4.759 4.885 4.955
SV-min 4.739 4.878 4.949
UNEDF1 4.740 4.887 4.958
@@
Appendix B Details of the shell model calculations and comparison with experiment
Here we provide further details of the shell model calculations, including valence spaces and potential truncations, and nuclear interactions. We also compare experimental nuclear observables to the theoretical shell model values, namely the nuclear energy levels, as well as the magnetic dipole and electric quadrupole moments and transitions. For each of these observables, we provide the values for 40Ar, 70,72,74,76Ge and 133Cs below, whilst the 73Ge and 127I comparisons can be found in reference [41].
Information on the ground state is most important for comparison with theoretical values, as it is used in the scattering calculations here. However, as the ground state of even-even nuclei is $0^{+}$, the magnetic dipole and electric quadrupole moment values are identically zero. Nevertheless, we employ the next non-zero state ($2_{1}^{+}$), which can give an indication of the overall accuracy of the calculation through information on the spectrum beyond the ground state.
40Ar
40Ar shell model calculations were performed in reference [41] with the $sdpf$ valence space, which constitutes single particle levels $1d_{5/2}$, $2s_{1/2}$, $1d_{3/2}$, $1f_{7/2}$, $2p_{3/2}$, $1f_{5/2}$ and $2p_{1/2}$, and assuming a 16O core. The valence space truncation employed consisted of an unrestricted $sd$ shell and an empty $pf$ shell for the protons, and a full $sd$ shell with an unrestricted $pf$ shell for the neutrons. Four shell model interactions were used: SDPF-NR [49], SDPF-U [50], EPQQM [51] and SDPF-MU [52].
The 40Ar theoretical energy levels are plotted against the experimental values in Fig. 5. The first two theoretical energy levels for all interactions are consistent with experiment within $\sim 200$ keV, except for SDPF-MU which is within $\sim 400$ keV from experiment for the $2_{1}^{+}$ state. Overall, the energy spectrum of the EPQQM interaction is least consistent with the other shell model calculations and with experiment, with the spectrum being much more sparse and spread out in the level spacings. Note that all calculations miss the first-excited 0+ state, which is due to a collective multiparticle-multihole excitation outside the valence space and will be remarked upon further in the following comparison with electromagnetic observables.
The magnetic dipole and electric quadrupole moments and transitions are presented in Table 8, where the effective $g$ factors and charges employed take on values of $e_{p}=1.5$, $e_{n}=0.5$, $g_{sp}=5.586$, $g_{sn}=-3.826$, $g_{lp}=1$, and $g_{ln}=0$. It is found that the shell model moments are not consistent with experimental values, and most experimental transitions are poorly reproduced as well.
Figure 5: Experimental energy levels (keV) for 40Ar are plotted up to 4000 keV and are compared with levels from the SDPF-NR, SDPF-U, SDPF-MU and EPQQM shell model calculations.
Table 8: Magnetic dipole and electric quadrupole moments, and electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions, between low-lying states of 40Ar. Experimental transition values are from [54], while moments are taken from [55, 56].
@@ Appendix B Details of the shell model calculations and comparison with experiment
70,72,73,74,76Ge
The germanium stable isotopes were studied in the unrestricted $f_{5}pg_{9}$ model space ($2p_{3/2}$, $1f_{5/2}$, $2p_{1/2}$ and $1g_{9/2}$ single-particle levels) with the JUN45 [57] and jj44b [58] shell model interactions in Ref. [41]. In addition, the GCN2850 interaction [59] was used in reference [38, 42] with a valence space truncation that consisted of limiting the occupation number of the $1g_{9/2}$ level to no more than two nucleons above the minimum occupation for all isotopes.
The experimental energy spectra are plotted against JUN45 and jj44b results for the even isotopes in Fig. 6. In the cases of 70,72,74Ge, the low-lying experimental spectra are a little better reproduced by the JUN45 interaction, which, unlike jj44b, was fitted to the Ge isotopes in consideration. The shell model spectra for 76Ge are practically identical to one another up to $\sim 1.9$ MeV. An agreement between theoretical and experimental levels of $\sim 200$ keV would be considered satisfactory for such shell model calculations using heavier nuclei, however this agreement exists only for some levels of the even isotopes. See also [41] for 73Ge results, where the theoretical spectra were found to be inconsistent with another, and not particularly consistent with experiment.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided for the even isotopes in Tables 9 and 10. For the magnetic dipole moments of the even isotopes, the experimental $2_{1}^{+}$ and $4_{1}^{+}$ states are well reproduced by both shell model interactions, whereas the jj44b value of the $2_{2}^{+}$ state is consistently closer to experiment compared to JUN45. Note that the compiler, Stone [55], has increased the uncetainties on the magnetic moments of the $2_{1}^{+}$ states reported by McCormick et al. [60] based on their statement that the absolute scale of the moments remains uncertain on the level of 20%. The $2_{1}^{+}$ electric quadrupole moment is quite inconsistent with the experimental values for 70,72,74Ge, where only the jj44b 76Ge value lies within the experimental range. Such discrepancies are not uncommon in medium-mass and heavy nuclei: the shell model is better equipped to describe single-nucleon features of the wavefunction than collective features like the shape of the charge distribution.
Figure 6: Experimental energy levels (keV) for germanium isotopes are plotted up to 2500 keV (70Ge), 2100 keV (72Ge), 2165 keV (74Ge) and 2200 keV (76Ge), and compared with levels from shell model calculations using the JUN45 and jj44b interactions.
Table 9: Magnetic dipole and electric quadrupole moments for low-lying states of 70,72,74,76Ge. Experimental moments are taken from [55, 56, 61, 60].
@@
Table 10: Electric quadrupole [B(E2)] transitions between low-lying states of 70,72,74,76Ge. Experimental transitions are taken from [62, 63, 64, 65].
@@
127I
Shell model calculations for 127I were performed in reference [41] in the model space which includes all proton and neutron orbits in the major shell between magic numbers 50 and 82, with single particle levels $1g_{7/2}$, $2d_{5/2}$, $2d_{3/2}$, $3s_{1/2}$ and $1h_{11/2}$, and with the SN100PN [66] and GCN5082 [59] interactions. The model space restrictions are described in reference [41], with the four different calculations given the names “SN100PN” and “GCN5082” (which share the same truncation), as well as “Truncated SN100PN” (with a more truncated scheme), and “GCN5082 Expanded” (with an alternative truncation to “GCN5082”). The valence space restriction employed in reference [38, 42] with the Baldridge and Dalton (B&D) interaction [67] involves fixing the occupation number of $1h_{11/2}$ to the minimum allowed nucleon number. See Ref. [41] for a comparison between shell model predictions and experimental nuclear structure data.
133Cs
Here, we consider the same valence space and shell model interactions (SN100PN and GCN5082) as 127I. We have performed four calculations for 133Cs, two of which employ the same truncations as the “SN100PN” and “GCN5082” calculations for 127I (and are given the same names here). In order to investigate the sensitivity of the shell model predictions to the truncation scheme, we consider two additional calculations, one each for the SN100PN and GCN5082 interactions, where the neutron $2d_{5/2}$ level now has a minimum of 4 neutrons (as opposed to previously being completely filled), with the remainder of the proton and neutrons levels retaining the previous truncations. This is to explore the effect of a less truncated valence space on the final results. These two additional calculations are termed “SN100PN Expanded” and “GCN5082 Expanded”.
The experimental energy spectra are plotted against the theoretical values for all the above calculations in Fig. 7. Only the SN100PN result reproduces the correct ground state spin and parity of $7/2^{+}$, however all four calculations reproduce the correct spins and parities of the three lowest-lying states within $200$ keV of experiment. For both shell model interactions, using a less truncated valence space causes the spacing of the levels to widen compared to the original calculations, generally pushing the spectrum further up. This is likely due to the opening up of the $2d_{5/2}$ level in these calculations, which increases pairing correlations in the ground state of the even-even configuration space that remains with the odd nucleon removed.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided in Tables 11 and 12. The proton and neutron effective charges and $g$ factors employed here are the same as those used in reference [41] for 127I. The experimental magnetic dipole moments are much better reproduced by the shell model calculations compared to the electric quadrupole ones. The experimental electric quadrupole transitions are only reproduced within the uncertainty margins by the GCN5082 calculation, specifically for transitions from the $5/2_{1}^{+}$ and $3/2_{1}^{+}$ states to the ground state. On the other hand, the theoretical magnetic dipole transition values are not consistent with experiment. However, the important test of the shell model wavefunction is the magnetic dipole moment, which is sensitive to the wavefunction of the single state, and to the single nucleon structure of the state, rather than to emerging collective properties, which affect the E2 transition rates and the quadrupole moments. The magnetic dipole transitions are suppressed among the dominant low-excitation configurations, as is evident from their small values, and so become very sensitive to small components in the wavefunction that are not of interest here. Despite the lack of agreement for some observables, the present shell model calculations can be trusted to include the important single-particle configurations in the ground-state wavefunction.
Figure 7: Experimental energy levels (keV) are plotted up to 750 keV for 133Cs, compared with levels from SN100PN and GCN5082 shell model calculations for two different valence space truncations (#2 indicates “Expanded” truncations).
Table 11: Magnetic dipole and electric quadrupole moments for low-lying states of 133Cs. Experimental moments are taken from [68, 55, 56, 69]. “#2” refers to the expanded truncation schemes.
@@
$5/2_{1}^{+}$66.40 -67.47 -63.36 -60.36 --22(5) [^1] From [69]. A different value, without sign, is given in Ref. [56]. 3.54 3.62 3.36 3.50 +3.45(2)
@@
Table 12: Electric quadrupole [B(E2)] and magnetic dipole [B(M1)] transitions between low-lying states of 133Cs. Experimental transition values are taken from [70]. “#2” refers to the expanded truncation schemes.
@@ Appendix C EFT/Nuclear Operators
The nuclear operators employed in Eq. (III.1) have the form
@@
The one-nucleon matrix elements in Eq. (III.1) have analytic expressions of the form [18]:
@@ Appendix D Form factors
In this section we plot the form factors for the individual responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ using the shell model results, where the $\Sigma^{\prime}$ response is only plotted for isotopes with non-zero nuclear ground state spin. This is presented for 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs, highlighting the corresponding COHERENT range of interest for each of these nuclei. The form factor expressions employed to plot these are presented in reference [41], except for 133Cs which is provided in Appendix D. The KN and SHF form factors were plotted using $(ZF\left(q,r_{\text{rms}}^{p})+NF(q,r_{\text{rms}}^{n})\right)^{2}$, while the shell model counterparts were plotted using $\sum_{N,N^{\prime}=p,n}F^{(N,N^{\prime})}_{X,X}(q^{2})$, where $F^{(N,N^{\prime})}_{X,X}(q^{2})$ is defined in Eq. (41).